OpenStudy (anonymous):

Can someone double check this for me please? let f(x93)=x^2+6x and g(x)=3-x. Find f(3)-g(3). f(3)-g(3)=27

OpenStudy (anonymous):

Is it f(x) or f(x = 93).. Correct it..

OpenStudy (anonymous):

f(3) = (3)^2 + 6(3) So, f(3) = 27 Now, g(3) = 3 - (3) = 0.. Just Subtract them and yes Your answer is very much correct...

OpenStudy (anonymous):

It's f(x), sorry, my typing is going down hill as the day drags on. Thank you very much.